OpenStudy (koikkara):

A 100 V voltmeter has full-scale accuracy of 5%. At its reading of 50 V it will give an error of? (8) A 100 V voltmeter has full-scale accuracy of 5%. At its reading of 50 V it will give an error of (a) 10% (b) 5% (c) 2.5% (d) 1.25% I got two solutions b and c, which one is right ? b ?

OpenStudy (anonymous):

Answer is a. If 100V is %100 then %5 is 5V Voltmeter can measure \[100\pm5V\] if 50V - %100 5V - %x x=100*5/50=%10 Sorry if answer is wrong. :)

OpenStudy (koikkara):

@Nursoltan Thank you for your valuable time ! As per the condition you have given the answer is right, but as per the question, A 100 V voltmeter has full-scale accuracy of 5%. means For a 100V= 5% then for 50V= ? leave it, consider how much u get for 1%=? I tried and found like... 1%= 100/5=20V (I'm not sure, is this right either) So, if 1%= 20 then 2%=40, 3%=60, 4%= 80 and 5%= 100 In this case, i get 2.5=50% Is this acceptable ?

OpenStudy (koikkara):

@.Sam. @Preetha Which answer looks correct ?

OpenStudy (koikkara):

Thank You in advance !